# Unisolver: PDE-Conditional Transformers Towards Universal Neural PDE Solvers

### v4

@@ 1 Introduction
Partial differential equations (PDEs) are essential for numerous scientific and engineering problems [8, 2], such as meteorology, electromagnetism and thermodynamics [41]. Since it is usually hard to obtain an analytic solution for a PDE, numerical methods are widely explored [1]. However, these numerical methods often require huge computation costs to generate a precise solution for each PDE. Recently, deep learning models have facilitated significant advancements across a wide range of domains [5, 24, 15] and have been applied to solving PDEs, i.e. neural PDE solvers [16]. Owing to their excellent capability to approximate nonlinear mappings, deep learning models can learn to fit pre-collected data [19] or physics-informed loss function [34] and generalize in a flash to new samples, providing an efficient approach to solving PDEs.
As shown in Figure 1, previous neural solvers can be broadly categorized into two paradigms: physics-informed neural networks (PINNs) [34] and neural operators [19]. The former trains deep models using a formalized PDE loss function, while the latter solely relies on pre-collected data. However, for PINNs, while formulating the PDE equations as objective functions ensures relatively accurate solutions, they struggle to generalize to new scenarios, necessitating retraining for each new task. Neural operators, on the other hand, directly learn from data and generalize better to diverse initial states and PDEs than PINNs. Nevertheless, purely based on training data may be insufficient to guide PDE solving. For example, in the case of a fluid governed by renowned Navier-Stokes equations, the typical task of neural operators is to predict future states based on past observations [19], while different viscosity coefficients and forcing terms will lead to distinct solutions even when the initial states stay the same. Thus, due to the omission of PDE information, current neural operators are mainly trained and tested on a limited set of PDEs. Notably, as neural solvers are expected to be efficient surrogate models of classical numerical solvers, generalization to various PDEs is essential for a practical neural solver.
@@
To tackle the generalization deficiency, several works have been proposed by incorporating the PDE information into deep models or training models with large-scale datasets. For example, message-passing neural PDE solver [3] concatenates PDE coefficients with the inputs. PDEformer [45] formalizes the PDE as a computation graph and employs the graph Transformer [46] to aggregate PDE information. Although these methods explore the potential of training models with both data and PDE information, they do not consider a complete set of PDE information, limiting their generalizability in some aspects. As for the other branches, such as DPOT [11], they purely scale up training sets with diverse PDEs and expect the generalizability emerges from large data and parameters. Although models can implicitly extract PDE information from observations, the extraction process is inherently complex and resembles the challenges associated with solving inverse problems [16]. Therefore, these models often end up fitting an insufficient or vague representation of the underlying observation distribution, which may ultimately hamper their generalizability to broader PDE solving.
@@ 2.1 Neural PDE Solvers
Previous neural PDE solvers can be roughly categorized into two paradigms [43]. The first is physics-informed neural networks (PINNs) [34], which optimize deep models by formalizing PDEs as objective functions. During training, model outputs and gradients gradually satisfy the targeted PDE, thereby instantiating the solution as a deep model. However, PINNs are usually hard to generalize to unseen PDEs, limiting their broader applications [41]. Another booming direction is neural operators, which learn from extensive data to approximate functional dependence between input and output Banach spaces [27, 18]. Among various neural operators, FNO [19] and its variants [21, 33, 42] are popular and well-established. FNO effectively approximates the kernel integral operator in the frequency domain. Recently, given the impressive progress achieved by Transformers [40], they have also been applied to solve PDEs. Existing methods treat inputs as sequences of tokens and adopt the attention mechanism to approximate integrals. GNOT [10] treats each mesh point as a token and utilizes the linear Transformer to address complexity issues. FactFormer [22] axially factorizes the attention block to boost efficiency. Transolver [43] learns intrinsic physical states as tokens behind input meshes, introducing a physics-attention mechanism. Despite the success of neural operators, they are only tested on the dataset with limited PDEs. Their effectiveness for various PDEs has not been fully explored.
@@ 2.2 Generalizable PDE Solvers
To guide the PDE-solving process, PDE information has been explored in many deep models. For example, PINO [20] imposes explicit equation constraints at a higher resolution to aid neural operator learning. CAPE [38] directly embeds PDE coefficients to adapt neural solvers to unseen equation coefficients. PROSE [23] and PITT [25] tokenize PDEs and embed mathematical expressions, informing the transformer of the underlying physics. PDEformer [45] represents symbolic equations as graphs and numeric components as nodes to optimize interactions between them. However, all of these methods, while incorporating equation information, do not leverage the mathematical structure of PDEs for complete and categorized embedding or integrating the prior information of equation symbols within the context of natural language. In contrast, Unisolver leverages the capabilities of large language models (LLMs) [39] to semantically embed the equation symbolic information and categorize the complete equation components based on mathematical insights for better PDE solving.
@@
As a vital cornerstone of deep learning [4, 12], recent research has also started to explore the effectiveness of large-scale training in solving PDEs. Subramanian et al. examine the scaling capabilities and transfer learning behaviors of FNO on three time-independent PDE families. MPP [29] proposes an auto-regressive strategy to train on a broad fluid mechanics-oriented benchmark. DPOT [11] enhances MPP with a denoising method and trains a Fourier Transformer on massive PDE data comprised of 12 datasets. PDEformer [45] focuses on a 1D time-dependent PDE family and pre-trains a graph transformer on 3M samples under various equation conditions. ICON [44] trains a single neural operator capable of performing in-context learning across a wide range of differential equations. However, most existing methods fall short of effectively and completely integrating PDE information, which Unisolver will address well.
@@ 3 Unisolver
To achieve ideal generalizability, we focus on the task of universal neural PDE solving shown in Figure 2. Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a bounded continuous set and =$\mathcal{M}=\{x_{1},\dots,x_{n}\}$ an $n$-point discretization of $\mathcal{D}$ recording coordinates of each point. For each observation pair, assume we have initial condition observations $\mathbf{X}$ as input and target quantities $\mathbf{Y}$ as output on the mesh $\mathcal{M}$, with governing PDE components $\mathcal{C}_{\text{PDE}}$ (e.g. PDE symbols, coefficients) that may vary for each observation. The universal neural PDE solving task is to approximate the input-PDE-output map $G\colon(\mathbf{X},\mathcal{M},\mathcal{C}_{\text{PDE}})\rightarrow\mathbf{Y}$ across a diverse training dataset and generalize in a flash to unseen PDEs.
@@ 3.1 Complete PDE Components
Here, we clarify the concept in the context of deep learning by considering the classical vibrating string equation with fixed endpoints as a motivating example, which can be solved explicitly, as shown in [8]. The analytical solution is provided in Appendix F.
@@
Domain geometry $[0,L]\times[0,T]$
@@ 3.2 Universal Components Embedding
As described in Section 3.1, PDE solutions are obtained by intricate interactions between initial conditions and complete equation components which can be grouped into two categories. In previous works [3, 38], these equation components are coarsely and incompletely included as conditions to modulate the input observations. In this paper, we will elaborate on how Unisolver finely and completely embeds all considered PDE components (Table 1) into deep PDE conditions based on our insights from the mathematical analysis.
@@
As shown in Table 1, other components can be categorized as domain-wise and point-wise based on their effect on the final solution. Correspondingly, we adopt different embedding methods for these two types. For domain-wise components, including coefficients represented as real-valued vectors and boundary types akin to class labels, we embed them using two linear layers with an intermediate SiLU activation function [7]. For point-wise components like external force, binary geometry mask, and boundary value functions are essentially physical fields observed on mesh $\mathcal{M}$. We apply the same patchify embedding method used for input observations, transforming them into deep representation sequences.
@@ 3.3 PDE-Conditional Transformer
We evenly split the hidden representations of the inputs along the channel dimension, with one half influenced by domain-wise deep conditions and the other half by point-wise deep conditions. In multi-head attention [40], our proposed subspace decoupling is equivalent to assigning some heads to learn the impact of domain-wise conditions while others focusing on point-wise conditions. This leads to improved representation learning for both categories, and minimized interference between deep PDE conditions from two categories.
@@
As shown in Figure 3, we aggregate conditions either before or after the attention and feedforward modules within Transformer. Inspired by recent conditional Transformers like DiT [31] and other conditional normalization approaches [30, 32], we take the aggregation paradigm to finely capture the intricate correlations between hidden inputs of initial observations and deep equation conditions. Specifically, we scale and shift the hidden representations of inputs based on the equation conditions. After passing through the Transformer modules, we use the equation conditions to softly select whether this information should be retained.
@@
Summarizing the above designs, we propose the Unisolver (Figure 3). Given input $\mathbf{X}$, it is projected to embeddings $\mathbf{X}^{0}$ using a patchify layer [6]. The complete PDE equation components $\mathcal{C}_{\text{PDE}}$ are embedded into deep conditions $\mathbf{C}_{\text{domain}}$ and $\mathbf{C}_{\text{point}}$ following Section 3.2. Suppose there are $N$ layers, the $n$-th layer of Unisolver can be formalized as:
@@
[…]
@@ 4 Experiments
As summarized in Table 2, three experimental large-scale benchmarks cover varied dimensions, resolutions and PDE components. The HeterNS is an extension of the NS dataset from FNO [2021a], incorporating multiple viscosity coefficients and external forces to enhance diversity. The 1D time-dependent PDEs, introduced by PDEformer [2024], is a large-scale dataset containing three million structured 1D PDE samples and evaluate the zero-shot generalization performance on PDEBench [2022]. The 2D mixed PDEs, collected by DPOT [2024], include 12 diverse datasets from four well-established benchmarks. More details can be found in Appendix G.
@@
We compare Unisolver with six advanced baselines on the HeterNS to demonstrate its generalizability under varied PDE components: the well-established FNO [2021a], PINO [2021b] and ViT [2020] and current state-of-the-art methods FactFormer [2023b], ICON [2023] and MPP [2023]. We augment these baselines by providing sufficient physics information to ensure a fair comparison, either by concatenating inputs with varied PDE components, providing prompting trajectories (ICON), or applying physics-informed loss (PINO).
Furthermore, we compare Unisolver with two generalizable solvers—PDEformer [2024] and DPOT [2024] on zero-shot generalization performance. We refrain from including additional baselines on these two benchmarks due to the substantial computational cost of using million-scale samples.
@@
All methods in the HeterNS benchmark are trained for 300 epochs using relative L2 loss and the ADAM optimizer [17] with an initial learning rate of $0.0005$ and a cosine annealing learning rate scheduler [26]. The batch size is set to 60. For the 1D time-dependent PDEs and 2D mixed PDEs, we follow the training strategies from the original papers of PDEformer [2024] and DPOT [2024] to ensure a fair comparison. Relative L2 is used as the evaluation metric. See Appendix H for full implementation details and hyper-parameter configurations of each model.
@@ 4.1 Heterogeneous 2D Navier Stokes Equation: HeterNS
We introduce HeterNS, an extension of the widely used 2D NS dataset [19], to assess how models handle diverse PDE components, particularly viscosity coefficients and force terms. It comprises five viscosity coefficients $\nu$ and three force terms differentiated by frequency $\omega$, resulting in 15 combinations of PDE components and 15,000 training samples. As depicted in Figure 4, we evaluate the model performance on in-distribution test with only unseen initial conditions and zero-shot generalization involving both unseen initial conditions and variations in viscosity coefficients or force terms.
@@
Table 5: Performance comparison (relative L2 (×$\times 10^{-2}$)) across 12 in-distribution test sets. We use a “source-PDE” format to denote different tasks (e.g. FNO-NS). The second row lists the primary PDE components for each dataset. See Appendix G.3 for details.
@@ Results
This benchmark contains three million high-quality 1D time-dependent PDE samples with varying equation formulations, coefficients, force terms and boundary conditions. We perform joint training on this extensive dataset, where the input for the training task includes all relevant PDE components, and the output records full space-time fields. After training, the model is evaluated across multiple test settings as depicted in Figure 5, including in-distribution test as well as zero-shot generalization on Burgers and Advection equations from PDEBench [37], which is another unseen dataset.
@@
This benchmark involves 12 datasets from four prominent benchmarks, covering a wide range of PDEs. After joint training on these diverse datasets, we perform in-distribution tests on each dataset as shown in Figure 6. Notably, the in-distribution test set also involves challenging variations in the PDE components. Moreover, unlike the balanced data in HeterNS, these datasets exhibit significant imbalances across different PDE components. To mitigate this issue, we adopt the balanced data sampling method from DPOT [11]; however, it still poses considerable challenges in managing such diverse PDE samples.
@@
As shown in Table 5, Unisolver outperforms DPOT [11] in 11 of 12 in-distribution test sets with an remarkable average promotion of 17.5% (5.50→ 4.54), except for the small Diffusion-Reaction (DR) dataset whose relative L2 tends to be saturated (less than 5%), verifying the effectiveness of our design. Unisolver shows consistently superior performance in several PDE component-dominated tasks, including coefficient generalization in FNO (2021a), force generalization in PDEArena (2023), and geometry generalization in CFDBench (2023), highlighting its ability to capture generalizable representations from complete PDE components.

### v5

@@ 1 Introduction
Partial differential equations (PDEs) are essential for numerous scientific and engineering problems (Evans, 2022; Arnol’d, 2013), such as meteorology, electromagnetism and thermodynamics (Wang et al., 2023). Since it is usually hard to obtain an analytic solution for a PDE, numerical methods are widely explored (Ames, 2014). However, these numerical methods often require huge computation costs to generate a precise solution for each PDE. Recently, deep learning models have facilitated significant advancements across a wide range of domains (Devlin et al., 2019; Liu et al., 2021; Jumper et al., 2021) and have been applied to solving PDEs, i.e. neural PDE solvers (Karniadakis et al., 2021). Owing to their excellent capability to approximate nonlinear mappings, deep learning models can learn to fit pre-collected data (Li et al., 2021a) or physics-informed loss function (Raissi et al., 2019) and generalize in a flash to new samples, providing an efficient approach to solving PDEs.
As shown in Figure 1, previous neural solvers can be broadly categorized into two paradigms: physics-informed neural networks (PINNs) (Raissi et al., 2019) and neural operators (Li et al., 2021a). The former trains deep models using a formalized PDE loss function, while the latter solely relies on pre-collected data. However, for PINNs, while formulating the PDE equations as objective functions ensures relatively accurate solutions, they struggle to generalize to new scenarios, necessitating retraining for each new task. Neural operators, on the other hand, directly learn from data and generalize better to diverse initial states and PDEs than PINNs. Nevertheless, purely based on training data may be insufficient to guide PDE solving. For example, in the case of a fluid governed by renowned Navier-Stokes equations, the typical task of neural operators is to predict future states based on past observations (Li et al., 2021a), while different viscosity coefficients and forcing terms will lead to distinct solutions even when the initial states stay the same. Thus, due to the omission of PDE information, current neural operators are mainly trained and tested on a limited set of PDEs. Notably, as neural solvers are expected to be efficient surrogate models of classical numerical solvers, generalization to various PDEs is essential for a practical neural solver.
@@
To tackle the generalization deficiency, several works have been proposed by incorporating the PDE information into deep models or training models with large-scale datasets. For example, message-passing neural PDE solver (Brandstetter et al., 2022) concatenates PDE coefficients with the inputs. PDEformer (Ye et al., 2024) formalizes the PDE as a computation graph and employs the graph Transformer (Ying et al., 2021) to aggregate PDE information. Although these methods explore the potential of training models with both data and PDE information, they do not consider a complete set of PDE information, limiting their generalizability in some aspects. As for the other branches, such as DPOT (Hao et al., 2024), they purely scale up training sets with diverse PDEs and expect the generalizability emerges from large data and parameters. Although models can implicitly extract PDE information from observations, the extraction process is inherently complex and resembles the challenges associated with solving inverse problems (Karniadakis et al., 2021). Therefore, these models often end up fitting an insufficient or vague representation of the underlying observation distribution, which may ultimately hamper their generalizability to broader PDE solving.
@@ 2.1 Neural PDE Solvers
Previous neural PDE solvers can be roughly categorized into two paradigms (Wu et al., 2024). The first is physics-informed neural networks (PINNs) (Raissi et al., 2019), which optimize deep models by formalizing PDEs as objective functions. During training, model outputs and gradients gradually satisfy the targeted PDE, thereby instantiating the solution as a deep model. However, PINNs are usually hard to generalize to unseen PDEs, limiting their broader applications (Wang et al., 2023). Another booming direction is neural operators, which learn from extensive data to approximate functional dependence between input and output Banach spaces (Lu et al., 2021; Kovachki et al., 2023). Among various neural operators, FNO (Li et al., 2021a) and its variants (Li et al., 2023a; Rahman et al., 2023; Wen et al., 2022) are popular and well-established. FNO effectively approximates the kernel integral operator in the frequency domain. Recently, given the impressive progress achieved by Transformers (Vaswani et al., 2017), they have also been applied to solve PDEs. Existing methods treat inputs as sequences of tokens and adopt the attention mechanism to approximate integrals. GNOT (Hao et al., 2023) treats each mesh point as a token and utilizes the linear Transformer to address complexity issues. FactFormer (Li et al., 2023b) axially factorizes the attention block to boost efficiency. Transolver (Wu et al., 2024) learns intrinsic physical states as tokens behind input meshes, introducing a physics-attention mechanism. Despite the success of neural operators, they are only tested on the dataset with limited PDEs. Their effectiveness for various PDEs has not been fully explored.
@@ 2.2 Generalizable PDE Solvers
To guide the PDE-solving process, PDE information has been explored in many deep models. For example, PINO (Li et al., 2021b) imposes explicit equation constraints at a higher resolution to aid neural operator learning. CAPE (Takamoto et al., 2023) directly embeds PDE coefficients to adapt neural solvers to unseen equation coefficients. PROSE (Liu et al., 2023) and PITT (Lorsung et al., 2024) tokenize PDEs and embed mathematical expressions, informing the transformer of the underlying physics. PDEformer (Ye et al., 2024) represents symbolic equations as graphs and numeric components as nodes to optimize interactions between them. However, all of these methods, while incorporating equation information, do not leverage the mathematical structure of PDEs for complete and categorized embedding or integrating the prior information of equation symbols within the context of natural language. In contrast, Unisolver leverages the capabilities of large language models (LLMs) (Touvron et al., 2023) to semantically embed the equation symbolic information and categorize the complete equation components based on mathematical insights for better PDE solving.
@@
As a vital cornerstone of deep learning (Brown et al., 2020; He et al., 2022), recent research has also started to explore the effectiveness of large-scale training in solving PDEs. Subramanian et al. examine the scaling capabilities and transfer learning behaviors of FNO on three time-independent PDE families. MPP (McCabe et al., 2023) proposes an auto-regressive strategy to train on a broad fluid mechanics-oriented benchmark. DPOT (Hao et al., 2024) enhances MPP with a denoising method and trains a Fourier Transformer on massive PDE data comprised of 12 datasets. PDEformer (Ye et al., 2024) focuses on a 1D time-dependent PDE family and pre-trains a graph transformer on 3M samples under various equation conditions. ICON (Yang et al., 2023) trains a single neural operator capable of performing in-context learning across a wide range of differential equations. However, most existing methods fall short of effectively and completely integrating PDE information, which Unisolver will address well.
@@ 3 Unisolver
To achieve ideal generalizability, we focus on the task of universal neural PDE solving shown in Figure 2. Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a bounded continuous set and $\mathcal{M}=\{x_{1},\dots,x_{n}\}$ an $n$-point discretization of $\mathcal{D}$ recording coordinates of each point. For each observation pair, assume we have initial condition observations $\mathbf{X}$ as input and target quantities $\mathbf{Y}$ as output on the mesh $\mathcal{M}$, with governing PDE components $\mathcal{C}_{\text{PDE}}$ (e.g. PDE symbols, coefficients) that may vary for each observation. The universal neural PDE solving task is to approximate the input-PDE-output map $G\colon(\mathbf{X},\mathcal{M},\mathcal{C}_{\text{PDE}})\rightarrow\mathbf{Y}$ across a diverse training dataset and generalize in a flash to unseen PDEs.
@@ 3.1 Complete PDE Components
Here, we clarify the concept in the context of deep learning by considering the classical vibrating string equation with fixed endpoints as a motivating example, which can be solved explicitly, as shown in (Evans, 2022). The analytical solution is provided in Appendix F.
@@
Domain geometry $0,L]\times[0,T]$
@@ 3.2 Universal Components Embedding
As described in Section 3.1, PDE solutions are obtained by intricate interactions between initial conditions and complete equation components which can be grouped into two categories. In previous works (Brandstetter et al., 2022; Takamoto et al., 2023), these equation components are coarsely and incompletely included as conditions to modulate the input observations. In this paper, we will elaborate on how Unisolver finely and completely embeds all considered PDE components (Table 1) into deep PDE conditions based on our insights from the mathematical analysis.
@@
As shown in Table 1, other components can be categorized as domain-wise and point-wise based on their effect on the final solution. Correspondingly, we adopt different embedding methods for these two types. For domain-wise components, including coefficients represented as real-valued vectors and boundary types akin to class labels, we embed them using two linear layers with an intermediate SiLU activation function (Elfwing et al., 2018). For point-wise components like external force, binary geometry mask, and boundary value functions are essentially physical fields observed on mesh $\mathcal{M}$. We apply the same patchify embedding method used for input observations, transforming them into deep representation sequences.
@@ 3.3 PDE-Conditional Transformer
We evenly split the hidden representations of the inputs along the channel dimension, with one half influenced by domain-wise deep conditions and the other half by point-wise deep conditions. In multi-head attention (Vaswani et al., 2017), our proposed subspace decoupling is equivalent to assigning some heads to learn the impact of domain-wise conditions while others focusing on point-wise conditions. This leads to improved representation learning for both categories, and minimized interference between deep PDE conditions from two categories.
@@
As shown in Figure 3, we aggregate conditions either before or after the attention and feedforward modules within Transformer. Inspired by recent conditional Transformers like DiT (Peebles & Xie, 2023) and other conditional normalization approaches (Park et al., 2019; Perez et al., 2018), we take the aggregation paradigm to finely capture the intricate correlations between hidden inputs of initial observations and deep equation conditions. Specifically, we scale and shift the hidden representations of inputs based on the equation conditions. After passing through the Transformer modules, we use the equation conditions to softly select whether this information should be retained.
@@
Summarizing the above designs, we propose the Unisolver (Figure 3). Given input $\mathbf{X}$, it is projected to embeddings $\mathbf{X}^{0}$ using a patchify layer (Dosovitskiy et al., 2020). The complete PDE equation components $\mathcal{C}_{\text{PDE}}$ are embedded into deep conditions $\mathbf{C}_{\text{domain}}$ and $\mathbf{C}_{\text{point}}$ following Section 3.2. Suppose there are $N$ layers, the $n$-th layer of Unisolver can be formalized as:
@@
where $n\in\{1,\ldots,N\}$, $\in\{\text{scale, shift, select}\}$, and ${\mathbf{X}}^{n}$ is the output of the $n$-th layer. Meanwhile, SA denotes self-attention, LN denotes layer normalization, and FF denotes feedforward layers in the Transformer model (Vaswani et al., 2017). Since the PDE components have a crucial impact on the range of the output, we scale and shift $\mathbf{X}^{N}$ based on the deep equation conditions, and then linearly project $\mathbf{X}^{N}$ to obtain the final output as predictions of $\mathbf{Y}$.
@@ 4 Experiments
As summarized in Table 2, three experimental large-scale benchmarks cover varied dimensions, resolutions and PDE components. The HeterNS is an extension of the NS dataset from FNO (2021a), incorporating multiple viscosity coefficients and external forces to enhance diversity. The 1D time-dependent PDEs, introduced by PDEformer (2024), is a large-scale dataset containing three million structured 1D PDE samples and evaluate the zero-shot generalization performance on PDEBench (2022). The 2D mixed PDEs, collected by DPOT (2024), include 12 diverse datasets from four well-established benchmarks. More details can be found in Appendix G.
@@
We compare Unisolver with six advanced baselines on the HeterNS to demonstrate its generalizability under varied PDE components: the well-established FNO (2021a), PINO (2021b) and ViT (2020) and current state-of-the-art methods FactFormer (2023b), ICON (2023) and MPP (2023). We augment these baselines by providing sufficient physics information to ensure a fair comparison, either by concatenating inputs with varied PDE components, providing prompting trajectories (ICON), or applying physics-informed loss (PINO).
Furthermore, we compare Unisolver with two generalizable solvers—PDEformer (2024) and DPOT (2024) on zero-shot generalization performance. We refrain from including additional baselines on these two benchmarks due to the substantial computational cost of using million-scale samples.
@@
All methods in the HeterNS benchmark are trained for 300 epochs using relative L2 loss and the ADAM optimizer (Kingma & Ba, 2015) with an initial learning rate of $0.0005$ and a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2016). The batch size is set to 60. For the 1D time-dependent PDEs and 2D mixed PDEs, we follow the training strategies from the original papers of PDEformer (2024) and DPOT (2024) to ensure a fair comparison. Relative L2 is used as the evaluation metric. See Appendix H for full implementation details and hyper-parameter configurations of each model.
@@ 4.1 Heterogeneous 2D Navier Stokes Equation: HeterNS
We introduce HeterNS, an extension of the widely used 2D NS dataset (Li et al., 2021a), to assess how models handle diverse PDE components, particularly viscosity coefficients and force terms. It comprises five viscosity coefficients $\nu$ and three force terms differentiated by frequency $\omega$, resulting in 15 combinations of PDE components and 15,000 training samples. As depicted in Figure 4, we evaluate the model performance on in-distribution test with only unseen initial conditions and zero-shot generalization involving both unseen initial conditions and variations in viscosity coefficients or force terms.
@@
Table 5: Performance comparison (relative L2 ($\times 10^{-2}$)) across 12 in-distribution test sets. We use a “source-PDE” format to denote different tasks (e.g. FNO-NS). The second row lists the primary PDE components for each dataset. See Appendix G.3 for details.
@@ Results
This benchmark contains three million high-quality 1D time-dependent PDE samples with varying equation formulations, coefficients, force terms and boundary conditions. We perform joint training on this extensive dataset, where the input for the training task includes all relevant PDE components, and the output records full space-time fields. After training, the model is evaluated across multiple test settings as depicted in Figure 5, including in-distribution test as well as zero-shot generalization on Burgers and Advection equations from PDEBench (Takamoto et al., 2022), which is another unseen dataset.
@@
This benchmark involves 12 datasets from four prominent benchmarks, covering a wide range of PDEs. After joint training on these diverse datasets, we perform in-distribution tests on each dataset as shown in Figure 6. Notably, the in-distribution test set also involves challenging variations in the PDE components. Moreover, unlike the balanced data in HeterNS, these datasets exhibit significant imbalances across different PDE components. To mitigate this issue, we adopt the balanced data sampling method from DPOT (Hao et al., 2024); however, it still poses considerable challenges in managing such diverse PDE samples.
@@
As shown in Table 5, Unisolver outperforms DPOT (Hao et al., 2024) in 11 of 12 in-distribution test sets with an remarkable average promotion of 17.5% (5.50→ 4.54), except for the small Diffusion-Reaction (DR) dataset whose relative L2 tends to be saturated (less than 5%), verifying the effectiveness of our design. Unisolver shows consistently superior performance in several PDE component-dominated tasks, including coefficient generalization in FNO (2021a), force generalization in PDEArena (2023), and geometry generalization in CFDBench (2023), highlighting its ability to capture generalizable representations from complete PDE components.